\begin{thebibliography}{10}

\bibitem{liu2017resource}
Zi-Wen Liu, Xueyuan Hu, and Seth Lloyd.
\newblock Resource destroying maps.
\newblock {\em Physical review letters}, 118(6):060502, 2017.

\bibitem{wigner1963information}
Eugene~P Wigner and Mutsuo~M Yanase.
\newblock Information contents of distributions.
\newblock {\em Proceedings of the National Academy of Sciences},
  49(6):910--918, 1963.

\bibitem{wigner1964positive}
Eugene~P Wigner and Mutsuo~M Yanase.
\newblock On the positive semidefinite nature of a certain matrix expression.
\newblock {\em Canad. J. Math}, 16:397--406, 1964.

\bibitem{luo2003wigner}
Shunlong Luo.
\newblock Wigner-yanase skew information and uncertainty relations.
\newblock {\em Physical review letters}, 91(18):180403, 2003.

\bibitem{luo2007skew}
Shunlong Luo and Qiang Zhang.
\newblock Skew information decreases under quantum measurements.
\newblock {\em Theoretical and Mathematical Physics}, 151(1):529--538, 2007.

\bibitem{luo2018coherence}
Shunlong Luo and Yuan Sun.
\newblock Coherence and complementarity in state-channel interaction.
\newblock {\em Physical Review A}, 98(1):012113, 2018.

\bibitem{chitambar2018quantum}
Eric Chitambar and Gilad Gour.
\newblock Quantum resource theories.
\newblock {\em arXiv preprint arXiv:1806.06107}, 2018.

\bibitem{horodecki2009quantum}
Ryszard Horodecki, Pawe{\l} Horodecki, Micha{\l} Horodecki, and Karol
  Horodecki.
\newblock Quantum entanglement.
\newblock {\em Reviews of modern physics}, 81(2):865, 2009.

\bibitem{winter2016operational}
Andreas Winter and Dong Yang.
\newblock Operational resource theory of coherence.
\newblock {\em Physical review letters}, 116(12):120404, 2016.

\bibitem{gour2008resource}
Gilad Gour and Robert~W Spekkens.
\newblock The resource theory of quantum reference frames: manipulations and
  monotones.
\newblock {\em New Journal of Physics}, 10(3):033023, 2008.

\bibitem{brandao2015reversible}
Fernando~GSL Brand{\~a}o and Gilad Gour.
\newblock Reversible framework for quantum resource theories.
\newblock {\em Physical review letters}, 115(7):070503, 2015.

\bibitem{coecke2016mathematical}
Bob Coecke, Tobias Fritz, and Robert~W Spekkens.
\newblock A mathematical theory of resources.
\newblock {\em Information and Computation}, 250:59--86, 2016.

\bibitem{chitambar2016comparison}
Eric Chitambar and Gilad Gour.
\newblock Comparison of incoherent operations and measures of coherence.
\newblock {\em Physical Review A}, 94(5):052336, 2016.

\bibitem{kollas2018optimization}
Nikolaos~K Kollas.
\newblock Optimization-free measures of quantum resources.
\newblock {\em Physical Review A}, 97(6):062344, 2018.

\bibitem{stormer2012positive}
Erling St{\o}rmer.
\newblock {\em Positive linear maps of operator algebras}.
\newblock Springer Science \& Business Media, 2012.

\bibitem{hiai2012quasi}
Fumio Hiai and D{\'e}nes Petz.
\newblock From quasi-entropy to various quantum information quantities.
\newblock {\em Publications of the Research Institute for Mathematical
  Sciences}, 48(3):525--542, 2012.

\bibitem{bhatia2013matrix}
Rajendra Bhatia.
\newblock {\em Matrix analysis}, volume 169.
\newblock Springer Science \& Business Media, 2013.

\bibitem{jenvcova2010unified}
Anna Jen{\v{c}}ov{\'a} and Mary~Beth Ruskai.
\newblock A unified treatment of convexity of relative entropy and related
  trace functions, with conditions for equality.
\newblock {\em Reviews in Mathematical Physics}, 22(09):1099--1121, 2010.

\bibitem{gour2017quantum}
Gilad Gour.
\newblock Quantum resource theories in the single-shot regime.
\newblock {\em Physical Review A}, 95(6):062314, 2017.

\bibitem{marvian2012symmetry}
Iman Marvian~Mashhad.
\newblock Symmetry, asymmetry and quantum information.
\newblock 2012.

\bibitem{marvian2013theory}
Iman Marvian and Robert~W Spekkens.
\newblock The theory of manipulations of pure state asymmetry: I. basic tools,
  equivalence classes and single copy transformations.
\newblock {\em New Journal of Physics}, 15(3), 2013.

\bibitem{streltsov2017colloquium}
Alexander Streltsov, Gerardo Adesso, and Martin~B Plenio.
\newblock Colloquium: quantum coherence as a resource.
\newblock {\em Reviews of Modern Physics}, 89(4):041003, 2017.

\bibitem{marvian2016quantify}
Iman Marvian and Robert~W Spekkens.
\newblock How to quantify coherence: Distinguishing speakable and unspeakable
  notions.
\newblock {\em Physical Review A}, 94(5):052324, 2016.

\bibitem{yu2017quantum}
Chang-shui Yu.
\newblock Quantum coherence via skew information and its polygamy.
\newblock {\em Physical Review A}, 95(4):042337, 2017.

\bibitem{vershynina2018quantum}
Anna Vershynina.
\newblock On quantum quasi-relative entropy.
\newblock {\em arXiv preprint arXiv:1810.00391}, 2018.

\bibitem{zhao2018coherence}
Haiqing Zhao and Chang-shui Yu.
\newblock Coherence measure in terms of the tsallis relative $\alpha$ entropy.
\newblock {\em Scientific reports}, 8(1):299, 2018.

\end{thebibliography}

\documentclass[%
 reprint,
 amsmath,amssymb,
 aps,
 pra,
]{revtex4-1}

\usepackage{graphicx}
\usepackage{dcolumn}
\usepackage{natbib}
\usepackage{bm}

\begin{document}
\title{Monotonicity of skew information and its applications in quantum resource theory}

\author{Weijing Li}%
 \email{liweijing13@mails.ucas.ac.cn}
\affiliation{%
Institute of Mathematics, Academy of Mathematics and Systems Science,\\
Chinese Academy of Sciences, Beijing, 100190, China
}%
\author{}
\affiliation{
 School of Mathematical Sciences,\\ University of Chinese Academy of Sciences, Beijing, 100049, China
}%
\author{}
\affiliation{
 UTS-AMSS Joint Research Laboratory for Quantum Computation and Quantum Information Processing,\\
 Academy of Mathematics and Systems Science, Chinese Academy of Sciences, Beijing, 100190, China
}%

\date{\today}

\begin{abstract}
We give an alternative proof of skew information via operator algebra approach and show its strong monotonicity under particular quantum TPCP maps. We then formulate a family of new resource measure if the resource can be characterized by a resource destroying map \cite{liu2017resource} and the free operation should be also modified. Our measure is easy-calculating and applicable to the coherence resource theory as well as quantum asymmetry theory. The operational interpretation needs to be further investigated.
\end{abstract}

\maketitle

\section{Introduction}
Wigner-Yanase skew information \cite{wigner1963information}\cite{wigner1964positive} was introduced originally to express the amount of information in density operator not commuting with the observables. For now there are a lot of generalizations and applications of skew information in quantum information sciences \cite{luo2003wigner}\cite{luo2007skew}\cite{luo2018coherence}. In the other hand, Quantum resource theories (QRTs) \cite{chitambar2018quantum} offer a highly versatile and powerful framework for studying different phenomena in quantum physics, typical examples are quantum entanglement \cite{horodecki2009quantum}, quantum coherence \cite{winter2016operational}, quantum reference frames \cite{gour2008resource} and Quantum Thermodynamics \cite{brandao2015reversible}. A general quantum resource theory consists of a class of ``free'' states along with a class
of ``free'' or allowable operations \cite{coecke2016mathematical}. The essential resource theoretic condition is that the set of free states is closed under the set of free operations. Hence, any state that is not free is a resource since it cannot be obtained using the allowable operations \cite{chitambar2016comparison}. In the quantum coherence theory, for example, there are various free operations such as incoherent operation (IO), dephasing- covariant operation (DIO) and strictly-incoherent operation (SIO). In conjunction with each of the operational classes, one
can define different measures of coherence. From a resource theoretic perspective, the crucial property of these measures is that they are monotonic under the specified class of operations. To give the measures physical meaning, one seeks to find some operational interpretation of the measure, thereby enabling
the measure to quantify some particular physical property or process. Various resource measures and resource monotones had been formulated and some operational interpretations had bee provided \cite{chitambar2018quantum} \cite{winter2016operational}.

Among various type of resources, one can specify the resource using the resource destroying maps \cite{liu2017resource}. Kollas Nikolaos showed a type of optimization-free measures based on resource destroying maps \cite{kollas2018optimization}.  Recently, Shunlong Luo $et\; al$ \cite{luo2018coherence} formulated a generalization of skew information \cite{luo2003wigner} from the viewpoint of state-channel interaction. So it is naturally to consider whether we can use this state-channel interaction to quantifying the resource given the resource destroying channel.

In this paper we first give an alternative proof of the monotonicity of skew information via operator algebra approach, and then prove that the skew information is of strong monotonicity, which is an open problem in \cite{luo2007skew}. Second we formulate a family of generalized skew information acted as proper resource measures in the framework of resource destroying map, the free operation should be also modified in the sense that it does not disturb the resource destroying map.

\section{Monotonicity of skew information}
Wigner and Yanase \cite{wigner1963information}\cite{wigner1964positive} introduced the notion of skew information of a density operator $\rho$ with respect to a self-adjoint observable $H$,
\begin{equation*}
    I(\rho,H)=-\frac{1}{2}\text{Tr}[\rho^{p},H][\rho^{1-p},H],
\end{equation*}
for $p=\frac{1}{2}$ and Dyson suggested extending this to $p\in (0,1)$. Recently Shunlong Luo \cite{luo2018coherence} gave some new generalizations. Given a Hilbert space $\mathcal{H}$, let $\mathcal{B}(\mathcal{H})$ denote the set of all operator on $\mathcal{H}$. For any operator $K \in \mathcal{B}(\mathcal{H})$, which needs not to be Hermitian, the skew skew information $I(\rho, K)$ of $\rho$ with respect to $K$ is defined as \cite{luo2018coherence}:
\begin{align*}
    I(\rho, K)= &||[\sqrt{\rho}, K]||^2= \text{Tr}[\sqrt{\rho}, K]^{\dagger}[\sqrt{\rho},K]\\
    =& \text{Tr}(\rho K^{\dagger}K)+\text{Tr}(\rho K K^{\dagger})-2\text{Tr}(\sqrt{\rho}K^{\dagger}\sqrt{\rho}K)
 \end{align*}
where $[X,Y]= XY-YX$ for all $X,Y \in \mathcal{B}(\mathcal{H})$ and $||X||^2= \text{Tr}(X^{\dagger}X)$ is the norm induced by the Hilbert-Schmidt inner product $\langle X, Y \rangle = \text{Tr} X^{\dagger}Y$. And if $\lambda$ is a Trace-Preserving and Completely Positive (TPCP) map with Kraus operators $K_i$, i.e., $\lambda(\rho)= \sum_i K_i \rho K_i^{\dagger}$, it is naturally to define the state-channel interaction \cite{luo2018coherence} by
\begin{equation*}
    I(\rho, \lambda)= \sum_i I(\rho, K_i).
\end{equation*}
The quantity $I(\rho, \lambda)$ enjoys some pleasant properties such as non-negativity, convexity, monotonicity, and $I(\rho, \lambda)$ is independent of the choice of Kraus operators of $\lambda$, see \cite{luo2018coherence} for more details. Among various properties of $I(\rho, \lambda)$, it is crucial that $I(\rho, \lambda)$ is monotone under some TPCP map $\mathcal{E}(\rho)=\sum_n M_n \rho M_n^{\dagger}$, i.e., 
\begin{equation*}
    I(\rho, \lambda) \ge I(\mathcal{E}(\rho), \lambda).
\end{equation*}
The original proof of monotonicity in \cite{luo2018coherence} is based on Landau-von Neumann equation. In this paper we give an alternative proof of the monotonicity of the skew information via the operator algebra method. And furthermore, we show that $I(\rho, \lambda)$ is of strong monotonicity, i.e., 
\begin{align*}
     I(\rho, \lambda) \ge &\sum_n p_n I(\rho_n, \lambda),
 \end{align*}
where $p_n= \text{Tr} M_n \rho M_n^{\dagger}$ and $\rho_n= M_n \rho M_n^{\dagger}/p_n $. To this end, we first recall some lemmas.

\textbf{Lemma 1} \cite{stormer2012positive}: If $\varphi$ is a unital positive map, then for every normal element $a$ in its domain, we have
$\varphi(a^{\dagger}a)\ge \varphi(a^\dagger)\varphi(a)$ and 
$\varphi(a^{\dagger}a)\ge \varphi(a) \varphi(a^\dagger)$.

\textbf{Lemma 2} \cite{hiai2012quasi}: Assume that $f: \mathbb{R}^+ \rightarrow \mathbb{R}$ is an operator monotone function with $f(0) \ge 0$ and $\alpha: \mathcal{B}(\mathcal{H}) \rightarrow \mathcal{B}(\mathcal{H})$ is a unital Schwarz mapping, that is, $\alpha(X^{\dagger}X) \ge \alpha(X^{\dagger})\alpha(X)$ for all $X \in \mathcal{B}(\mathcal{H})$, then
\begin{equation*}
    S^A_f(\alpha^{\dagger}(\rho_1) || \alpha^{\dagger}(\rho_2)) \ge S_f^{\alpha(A)}(\rho_1, \rho_2)
\end{equation*}
for $A \in \mathcal{B}(\mathcal{H})$ and for invertible density operators $\rho_1, \rho_2 \in \mathcal{B}(\mathcal{H})$, where 
\begin{align*}
    &S_f^A(\rho_1 || \rho_2)\\
    =&\langle A \rho_2^{1/2}, f(\Delta(\rho_1/\rho_2))(A \rho_2^{1/2})\\
    =&\text{Tr} \rho_2^{1/2}A^{\dagger}f(\Delta(\rho_1/\rho_2))(A \rho_2^{1/2}),
\end{align*}
and $\Delta(\rho_1 /\rho_2): \mathcal{B}(\mathcal{H}) \rightarrow \mathcal{B}(\mathcal{H})$ is the linear mapping defined by
\begin{equation*}
    \Delta(\rho_1/\rho_2)(X)=\rho_1 X \rho_2^{-1}.
\end{equation*}

\textbf{Lemma 3}: For a TPCP map $\mathcal{E}$ acted as $\mathcal{E}(\rho)=\sum_n M_n \rho M_n^{\dagger}$ and an operator $K \in \mathcal{B}(\mathcal{H})$,
$[K, M_n]=0$ if and only if 
$\mathcal{E}^{\dagger}(K)=K$ and $\mathcal{E}^{\dagger}(K^{\dagger}K)=K^{\dagger}K$; 
$[K^{\dagger}, M_n]=0$ if and only if 
$\mathcal{E}^{\dagger}(K)=K$ and $\mathcal{E}^{\dagger}(KK^{\dagger})=KK^{\dagger}$, where $\mathcal{E}^{\dagger}$ is the adjoint of $\mathcal{E}$ in the sense that $\langle X, \mathcal{E}(Y)\rangle = \langle \mathcal{E}^{\dagger}(X), Y\rangle$.

\textit{Proof.}---The necessity is very simple since if $[M_n, K]=0$ for all $n$ then $M_n K=K M_n$ and then $K^{\dagger}M_n^{\dagger}=M_n^{\dagger}K^{\dagger}$, the conclusion follows.

 For the sufficiency we only need to consider the equality
\begin{align*}
     &\sum_n [M_n, K]^{\dagger}[M_n, K]\\
    =&K^{\dagger}K-\mathcal{E}^{\dagger}(K^{\dagger})K-K^{\dagger}\mathcal{E}^{\dagger}(K)+\mathcal{E}^{\dagger}(K^{\dagger}K).
\end{align*}
The second statement is followed similarly.

We then give an proof of the monotonicity as well as strong monotonicity of $I(\rho, \lambda)$.

\textbf{Theorem 4}: For a TPCP map $\lambda(\rho)= \sum_i K_i \rho K_i^{\dagger}$, $I(\rho, \lambda)$ is well defined:
\begin{equation*}
    I(\rho, \lambda)= \sum_i I(\rho, K_i).
\end{equation*}
If TPCP map $\mathcal{E}(\rho)=\sum_n M_n \rho M_n^{\dagger}$ satisfies $\mathcal{E}^{\dagger}(K_i)= K_i$ and $\mathcal{E}^{\dagger}(K_i^{\dagger}K_i)=K_i^{\dagger}K_i$ as well as $\mathcal{E}^{\dagger}(K_i K_i^{\dagger})=K_i K_i^{\dagger}$, then $I(\rho, \lambda)$ is monotone and strong monotone under the action of $\mathcal{E}$:
 \begin{align*}
     I(\rho, \lambda) \ge &I(\mathcal{E}(\rho), \lambda);\\
     I(\rho, \lambda) \ge &\sum_n p_n I(\rho_n, \lambda),
 \end{align*}
where $p_n= \text{Tr} M_n \rho M_n^{\dagger}$ and $\rho_n= M_n \rho M_n^{\dagger}/p_n $.
 
 \textit{Proof.}---
 Since $\mathcal{E}$ is a TPCP map, then by Lemma 1  $\mathcal{E}^{\dagger}$ is a unital completely positive map and satisfies the Schwarz inequality: $\mathcal{E}(X^{\dagger}X) \ge \mathcal{E}(X^{\dagger}) \mathcal{E}(X)$.
 For a fixed Kraus operator $K_i$ of $\lambda$, $\mathcal{E}^{\dagger}(K_i)=K_i$ and $\rho = \rho_1 = \rho_2$, consider the function $f(x)=\sqrt{x}$, it is easy to see that $f(x)$ is operator concave as well as operator monotone \cite{bhatia2013matrix}, by Lemma 2, we have
 \begin{equation*}
     \text{Tr}(K_i \sqrt{\rho} K_i^{\dagger} \sqrt{\rho}) \le \text{Tr}(K_i \sqrt{\mathcal{E}(\rho)} K_i^{\dagger} \sqrt{\mathcal{E}(\rho)}).
 \end{equation*}
 Since $\mathcal{E}^{\dagger}(K_i^{\dagger}K_i)=K_i^{\dagger}K_i$ and $\mathcal{E}^{\dagger}(K_i K_i^{\dagger})=K_i K_i^{\dagger}$, by taking inner product with respect to $\rho$, we have
 \begin{align*}
     \text{Tr}(K_i \rho K_i^{\dagger})=&\text{Tr}(K_i \mathcal{E}(\rho) K_i^{\dagger}),\\
     \text{Tr}(\rho K_i K_i^{\dagger})=&\text{Tr}(\mathcal{E}(\rho)K_i K_i^{\dagger}).\\
 \end{align*}
 Since these two terms are both linear on $\rho$, we can conclude that
 \begin{align*}
     &I(\rho, K_i)\\
     =&\text{Tr}( \rho K_i^{\dagger}K_i)+\text{Tr}(\rho K_i K_i^{\dagger})-2\text{Tr}(K_i \sqrt{\rho} K_i^{\dagger} \sqrt{\rho})\\
     \ge &I(\mathcal{E}(\rho), K_i),
 \end{align*}
 by summing up all $i$ it must holds that 
 \begin{align*}
     I(\rho, \lambda) \ge I(\mathcal{E}(\rho), \lambda).
 \end{align*}
 
 Thus we reproduce the monotonicity result of $I(\rho, \lambda)$ in \cite{luo2018coherence} via the operator algebra viewpoint. Next we consider the strong monotonicity of $I(\rho, \lambda)$.
 
 For the free operation $\mathcal{E}(\rho)=\sum_n M_n \rho M_n^{\dagger}$, consider the channel
 \begin{align*}
      &\tilde{\mathcal{E}}(\rho \otimes |0\rangle \langle 0|)\\
     =&\sum_n M_n \otimes U_n (\rho \otimes |0\rangle \langle 0|) M_n^{\dagger}\otimes U_n^{\dagger}\\
     =&\sum_n M_n \rho M_n^{\dagger} \otimes |n\rangle \langle n|,
 \end{align*}
where $U_n$ is a unitary operator $\sum_{n=0}^{N-1} |k+n \; (\text{mod}\, N)\rangle \langle k|$ and $N$ is the number of Kraus operators $M_n$. It is easy to verify that
\begin{align*}
    \tilde{\mathcal{E}}^{\dagger}(K_i\otimes I)=K_i \otimes I
\end{align*}
 and \begin{align*}
     \tilde{\mathcal{E}}^{\dagger}(K_i^{\dagger}K_i\otimes I)=&K_i^{\dagger}K_i\otimes I,\\
     \tilde{\mathcal{E}}^{\dagger}(K_i K_i^{\dagger}\otimes I)=&K_i K_i^{\dagger}\otimes I,
 \end{align*}
 by the monotonicity of $I(\rho, \lambda)$, therefore,
 \begin{align*}
    &I(\rho, \lambda)\\
    =&I(\rho \otimes |0\rangle \langle 0|, \lambda \otimes I)\\
    \ge&I(\sum_n M_n \rho M_n^{\dagger} \otimes |n\rangle\langle n|, \lambda \otimes I)\\
    =&\sum_n p_n I(\rho_n, \lambda).
 \end{align*}
This finished the proof.

In \cite{luo2007skew}, the authors showed that for Hermitian operator $H$, $I(\rho,H)$ is monotone under a TPCP map $\mathcal{E}(\rho)=\sum_n M_n \rho M_n^{\dagger}$:
\begin{equation*}
    I(\rho, H) \ge I(\mathcal{E}(\rho),H)
\end{equation*}
as long as $\mathcal{E}$ satisfies: $\mathcal{E}^{\dagger}(H)=H$ and $\mathcal{E}^{\dagger}(H^2)=H^2$, or equivalently, $[H,M_n]=0$ for all $n$, but they only proved
that $I(\rho, H)$ satisfies the strong monotonicity in two-dimensional case. In fact, by same line reasoning in Theorem
4, we can show that $I(\rho, H)$ satisfies the strong monotonicity in general. Virtually,
\begin{align*}
    &I(\rho, H)\\
    =&I(\rho \otimes |0\rangle \langle 0|, H \otimes I)\\
    \ge&I(\tilde{\mathcal{E}}(\rho \otimes |0\rangle \langle 0|, H \otimes I))\\
    =&I(\sum_n M_n \rho M_n^{\dagger}\otimes |n\rangle \langle n|, H\otimes I)\\
    =&\sum_n p_n I(\rho_n, H).
\end{align*}
Hence we solve the open problem leaved in \cite{luo2007skew}.

We can even do further. In \cite{jenvcova2010unified} the authors formulated a family of functions 
\begin{equation*}
    g_p(x)=\begin{cases}
        \frac{1}{p(1-p)}(x-x^p) \quad &p \ne 1\\
        x \log x \quad &p=1, 
    \end{cases}
\end{equation*}
where $x>0$ and $p \in (0,2]$. For strictly positive $A,B$ they define \cite{jenvcova2010unified}
\begin{align*}
    &J_p(K, A, B)\\
    \equiv & \left<(K\sqrt{B}),  g_p(\Delta(A/B))(K\sqrt{B})\right>\\
    =&\begin{cases}
    \frac{1}{p(1-p)}(\text{Tr}K^{\dagger}AK - \text{Tr}(K^{\dagger}A^p K B^{1-p}))\;\\\quad \quad \quad \quad \quad \quad \quad \quad \quad 
    p\in(0,1)\cup (1,2),\\
    \text{Tr}KK^{\dagger}A \log A- \text{Tr}K^{\dagger}AK\log B \quad p=1,\\
    -\frac{1}{2}(\text{Tr}K^{\dagger}AK-\text{Tr}AKB^{-1}K^{\dagger}A) \quad p=2.
    \end{cases}
\end{align*}
One can see that this definition generalizes the relative entropy as well as the skew information. In fact, when $p=1$ and $K=I$, $J_p(K,A,B)$ reduces to the usual relative entropy, i.e., 
\begin{equation*}
    J_1(I,A,B)=\text{Tr}(A\log A-A\log B).
\end{equation*}

And when $p=\frac{1}{2}$, $K=K^{\dagger}$ and $A=B$,
\begin{equation*}
    J_{\frac{1}{2}}(K,A,A)=-\frac{1}{2p(1-p)}\text{Tr}[K,A^p][K,A^{1-p}],
\end{equation*}
 which yields the original skew information up to a constant.
 
 Since $J_p(K,A,B)$ enjoys some pleasant properties like skew information, it is naturally to extend our previous discussion by utilizing this unified entropy.
 
 Define 
 \begin{equation*}
    I_p(\rho, K)= J_p(K, \rho, \rho) 
 \end{equation*} for operator $K$ and density operator $\rho$ and $p\in [\frac{1}{2},1]$, and when a TPCP map $\lambda$ written as $\lambda(\rho)=\sum_i K_i \rho K_i^{\dagger}$, we define
 \begin{equation*}
     I_p(\rho, \lambda)=\sum_i I_p(\rho, K_i)
 \end{equation*}
 for $p\in [\frac{1}{2},1]$. If $p=\frac{1}{2}$,  $I_p(\rho,\lambda)$ is very similar to $I(\rho, \lambda)$ in Theorem 4 but there are several subtleties due to the linear term. As a consequence, we only need to require that $\mathcal{E}^{\dagger}(K K^{\dagger})=K K^{\dagger}$, or equivalently, $[K, M_n^{\dagger}]=0$ for all $n$. We must show that $I_p(\rho, \lambda)$ share the same properties as $I(\rho, \lambda)$ when $p\in [\frac{1}{2},1]$. We first show that $I_p(\rho, \lambda)$ is independent of the Kraus operator representations and thus $I_p(\rho, \lambda)$ is indeed a quantity about $\rho$ and channel $\lambda$. In fact, if $\lambda(\rho)=\sum_i E_i \rho E_i^{\dagger}=\sum_j K_j \rho K_j^{\dagger}$ for all density operator $\rho$, then there exists a unitary matrix $U=(u_{ij})$ such that $E_i=\sum_j u_{ij}K_j$. By some direct calculation one can see that $I_p(\rho, \lambda)=\sum_i I_p(\rho, E_i)=\sum_j I_p(\rho, K_j)$, which is independent of the Kraus operators of $\lambda$. The non-negativity and convexity of $I_p(\rho,\lambda)$ can be found in \cite{jenvcova2010unified}. For $p\in (0,1)$, the function $f(x)=x^p$ is operator monotone as well as operator concave \cite{bhatia2013matrix}. So the monotonicity and strong monotonicity of $I_p(\rho, \lambda)$ under TPCP map $\mathcal{E}(\rho)=\sum_n M_n \rho M_n^{\dagger}$ follows from the same line in Theorem 4. However, there does not exist any discussion about the monotonicity when $1 \le p \le 2$. Since in this case, the function $f(x)=x^p$ is a operator convex function, the method in Lemma 2 is not available anymore. We leave it as an open problem. For now we can conclude the following theorem.

 \textbf{Theorem 5}: For a TPCP map $\lambda(\rho)= \sum_i K_i \rho K_i^{\dagger}$ with Kraus operators $K_i$, the other TPCP map $\mathcal{E}(\rho)=\sum_n M_n \rho M_n^{\dagger}$ satisfies $\mathcal{E}^{\dagger}(K_i)= K_i$ and $\mathcal{E}^{\dagger}(K_i K_i^{\dagger})=K_i K_i^{\dagger}$ for all $i$ and $p\in (0,1)$, the quantity 
\begin{equation*}
    I_p(\rho, \lambda)= \sum_i I_p(\rho, K_i)
\end{equation*}
 is nonnegative, convex on $\rho$ and monotonicity under the action of $\mathcal{E}$. That is, $I(\rho, \lambda) \ge 0$ with equality if and only if $\lambda(\rho)=\rho$, $I(\rho, \lambda)$ is convex on $\rho$ and more crucially,
 \begin{equation*}
     I_p(\rho, \lambda) \ge I_p(\mathcal{E}(\rho), \lambda).
 \end{equation*}
Furthermore, it is of strong monotonicity:
\begin{equation*}
     I_p(\rho, \lambda) \ge \sum_n p_n I_p(\rho_n, \lambda),
\end{equation*}
where $p_n= \text{Tr} M_n \rho M_n^{\dagger}$ and $\rho_n= M_n \rho M_n^{\dagger}/p_n $.

\section{Applications in Resource framework}
Given a Hilbert space $\mathcal{H}$, we say that a TPCP map $\lambda: \mathcal{B}(\mathcal{H}) \rightarrow \mathcal{B}(\mathcal{H})$ is a resource destroying map \cite{liu2017resource} if it satisfies the following two properties:
\begin{enumerate}
    \item It maps any free state $\rho \in \mathcal{F}(\mathcal{H})$ to itself; i.e., $\lambda(\rho)=\rho$.
    \item It maps any (possibly not free) density operator $\rho$ to a free state; i.e., $\lambda(\rho) \in \mathcal{F}(\mathcal{H})$.
\end{enumerate}
From its definition, it is not clear at all that a resource destroying map exists for a given QRT. However, the full necessary and sufficient conditions for the existence of a resource destroying map were derived in \cite{gour2017quantum}.

In this section we show that the skew information $I(\rho, \lambda)$ and its generalization $I_p(\rho, \lambda)$ can be acted as a proper resource measure with respected to resource destroying map $\lambda$. The typical example is in quantum asymmetry theory. The resource theory of asymmetry with respect to a given representation of a symmetry group $G$ has been used extensively to distinguish and quantify the symmetry-breaking properties of both the states and the operations \cite{marvian2012symmetry} \cite{marvian2013theory}. The asymmetry theory can be described by a resource destroying map \cite{liu2017resource}, namely,
\begin{equation*}
    \lambda(\rho)=\int d\mu U_g \rho U_g^{\dagger},
\end{equation*}
where $U_g$ is the unitary representation of $g\in G$ and $d\mu$ is the Haar measure with respect to $G$. For simplicity we confine our focus on finite group $G$ but note that the conclusions hold also for compact Lie group. The symmetric state of the free state are those invariant under $\lambda$, i.e., 
\begin{equation*}
    \mathcal{F}=\{\rho \;| \lambda(\rho)=\rho\}.
\end{equation*}
An equivalently characterization is that \cite{marvian2012symmetry}
\begin{equation*}
    \mathcal{F}=\{\rho\;| \;U_g \rho U_g^{\dagger}=\rho\; \text{for all}\; g\in G\}.
\end{equation*}
The free operation we consider in this paper is the TPCP map $\mathcal{E}(\rho)=\sum_n M_n \rho M_n^{\dagger}$ such that $\mathcal{E}^{\dagger}(U_g)=U_g$ for all $g \in G$. Since $U_g$ are unitary thus normal the second type of  condition $\mathcal{E}^{\dagger}(U_g U_g^{\dagger})=U_g U_g^{\dagger}=I$ is satisfied automatically. By Lemma 3, it holds that $[M_n, U_g]=0$ for all $n$ and all $g \in G$. Therefore, the generalized skew information between state and channel $I_p(\rho, \lambda)$ can be served as a proper resource measure of asymmetry due to Theorem 5.

Another typical example is the resource theory of coherence \cite{winter2016operational} \cite{streltsov2017colloquium} \cite{marvian2016quantify} \cite{yu2017quantum}. In this case the resource destroying map is the dephasing map \begin{equation*}
    \Delta(\rho)=\sum_i |i\rangle \langle i|\rho|i\rangle \langle i|
\end{equation*} 
for a fixed basis $\{|i\rangle\}$. The Kraus operators are $K_i=|i\rangle \langle i|$ and for $p=\frac{1}{2}$, $I_\frac{1}{2}(\rho, \Delta)$ reproduces the result in \cite{yu2017quantum} up to a constant.

\section{Discussions}
In this paper we formulated a family of resource measure extended by skew information when the resource theory can be characterized by a resource destroying map and for a class of particular free operations. The new measure acted well in asymmetry theory and coherence theory. There are also some unsolved problems. The first one is how to generalize our measure to a more wide free operations. After all, the requirement for commute relation between Kraus operator seems too severe. The second problem is to consider the monotonicity of $I_p(\rho, \lambda)$ under the TPCP maps when $p \in (1,2]$. Recently, the author in \cite{vershynina2018quantum} showed the monotonicity of quantum quasi-entropy under partial trace, but it is different from our open problem. It is expected that a similar operator algebra approach proof should be applied. And the last one is to figure out the relationship between our measure and the measure introduced in \cite{zhao2018coherence}.

\section{Acknowledgments}
The author is very grateful to professor Shunlong Luo, professor Shaoming Fei
and Wei Xie for insightful discussions.
\bibliographystyle{unsrt}
\bibliography{bib.bib}
\end{document}